Mortgage

You plan to purchase an $130,000 house using a 30-year mortgage obtained from your local bank. The mortgage rate offered to you is 4.75 percent. You will make a down payment of 10 percent of the purchase price.
 
a. Calculate your monthly payments on this mortgage.
b. Calculate the amount of interest and, separately, principal paid in the 240th payment.
c. Calculate the amount of interest and, separately, principal paid in the 270th payment.
d. Calculate the amount of interest paid over the life of this mortgage.
Answer #1

Value of House = 130000, Down Payment = 10%

So, Loan Amount = 130000*0.9 = 117000

Mortgage Rate = 4.75%. So, monthly rate = 4.75%/12 = 0.3958%

Part a

Monthly payment can be calculated using PV of annuity formula:

PV = C*((1-1/(1+r)T)/r), where C = Installment, r = Periodic Interest Rate, T = number of periods

Here, T =30*12 = 360, r = 0.3958%, PV = 117000

So, C =117000/((1-1/(1+0.3958%)360)/0.3958%) = 610.33

Part b

Amount of Interest = $231.92

Amount of principal = $378.41

Part c

Amount of interest = $184.30

Amount of principal = $426.03

Part d

Total Amount paid over the life or mortgage = 610.33*360 = 219718.86

Principal = 117000

So, Amount of Interest Paid = 219718.8 - 117000 = $102717.86

*Answer for part b & c was derived from the amortization schedule

Here is a snapshot of the amortization schedule:

1 Loan Amount 2 Interest Rate 117000 4.7500% annual 0.3958% Month! 4 PrincipalPrincipal Repaid Remaining 5 Month Principal In

You need to fill in the month column with values 1 to 360. After that all you need to do is do the calculation for month 1 and month 2 as per the formula used (highlighted in yellow). Then select the cells corresponding month 2 and drag it down to month 360. Everything will get auto-calculated in excel.